Electron scattering by isolated atoms and molecules occurs in the gas phase. It plays a key role in plasma physics and chemistry and it's important for such applications as semiconductor physics.




Electron scattering experiment provides most accurate values of the sizes of atomic nuclei.




Estimate the rough energy of electron required for this.




Ans: ~250 MeV

1
Expert's answer
2022-03-23T16:14:03-0400

E = p2/(2 * m) = h2/(2 * m * r2) = (h * c)2/(2 * m * r2 * c2) = (197.3MeVF)2/(2*(940MeV)*(4F)2)